# Surface-Aware Distilled 3D Semantic Features

### v1

@@ 4.1 Setup
Unlike other approaches [41, 21], our method does not require canonical mesh alignment [41], nor rely on its estimation [21] because it purely uses intrinsic mesh properties rather than absolute vertex position. Similar considerations were made by prior methods based on Functional maps [58]. Furthermore, our encoder is point-based and does not take into account global shape and connectivity. Both of these design choices favor generalization under transformations ranging from coordinate swap to shape reposing.
4.2 Separating Front Paw from Back Paw
Our embedding aims to separate multiple instances of the same class that are difficult to directly disambiguate in $\mathcal{F}_{\mathcal{M}}$. For example, consider the two surface points, $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, on the bear’s paws in Fig. 2. The prevalent semantic significance of the “paw” concept hinders the separability of the corresponding base features $\mathbf{f}_{1}$ and $\mathbf{f}_{2}$. Fig. 6 illustrates this for human arms and Diff3F features [24]. To solve this, we train a point-wise feature autoencoder, producing our surface-aware features $\mathcal{S}_{\mathcal{M}}\subset\mathbb{R}^{s}$ in its embedding space. We motivate the feasibility of separation by the prior observations that vision features additionally carry information about the global pose [91]. We postulate that this enables our model to distinguish between part instances when guided by their intrinsic distance. We adopt the geodesic distance $d_{1,2}$ between $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ for this purpose. Following contrastive learning, we sample point pairs to enforce $\phi(\mathbf{s}_{1},\mathbf{s}_{2})\approx d_{1,2}$ for $\mathbf{s}_{n}\in\mathcal{S}_{\mathcal{M}}$. We validate our choice of hyperspherical embedding space against Euclidean space in Sec. 5.3.
@@ 4.2 Separating Front Paw from Back Paw
We train a base feature encoder $\mathcal{E}$, such that $\mathbf{s}_{n}=\mathcal{E}(\mathbf{f}_{n})/\lVert\mathcal{E}(\mathbf{f}_{n})\lVert_{2}$ is surface-aware feature $\mathbf{s}_{n}\in\mathbb{R}^{s}$ in a hypersphere embedding. During training, we randomly sample a training mesh $\mathcal{M}\in\mathbf{M}_{t}$ with base features $\mathcal{F}_{\mathcal{M}}$, which we encode pointwise to obtain $\mathcal{S}_{\mathcal{M}}$.
In each training iteration, we use furthest-point sampling to choose a random subset of $A$ anchor points $\mathbf{p}_{a}$ among the mesh vertices $\mathbf{p}_{i}\in\mathcal{M}{}$ and compute geodesic distances $d_{n,a}$ for each pair of a mesh and anchor point. We additionally rescale $d_{n,a}$ to a maximum of one, such that $d^{\prime}_{n,a}:=d_{n,a}/\max_{n,a}(d_{n,a})$, which removes the dependency on the scale of the mesh.
From here, our contrastive loss enforces preservation of the geodesic distances in the embedding space:
$\mathcal{L}_{c}=\frac{1}{NA}\sum_{n}^{N}\sum_{a}^{A}|d^{\prime}_{n,a}-\left(1-\frac{\phi(\mathbf{s}_{n},\mathbf{s}_{a})}{2}\right)|.$ (2)
The loss operates in a hyperspherical embedding space and utilizes cosine similarity mapped to the $[0,1]$ range. Hereby, $\mathcal{L}_{c}$ penalizes features close in the embedding space but distant on the shape surface and vice versa.
@@
$\mathcal{L}_{r}=\frac{1}{N}\sum_{n}1-\phi(\mathbf{f}_{n},\mathbf{\bar{f}}_{n}).$ (3)
@@
Note that our training procedure, in contrast to related works [44, 21, 14], does not require target and source pairs.
4.3 Implementation
@@
During preprocessing, we rasterize our triangular meshes and precompute base features for all vertices following Sec. 3. We implement our autoencoder in PyTorch2 [3]. The encoder $\mathcal{E}$ is a Multilayer Perceptron consisting of three blocks, where each block has two linear layers, SiLU activation [29], and layer normalization [6]. The first layer in each block employs a skip connection [39], while the second reduces the dimensionality by a factor of two. For Diff3F [24] as base features, the dimensionality is reduced from $f=2048$ to $s=256$. The decoder $\mathcal{D}$ is a mirrored copy of the encoder. We train our model on NVIDIA RTX 3090 for 50k iterations with the AdamW optimizer [49] and a learning rate of $0.0001$ that take $\approx 2$ hours.
We choose an exponential moving average [60] of the model with the lowest validation loss, without the need for any correspondence labels. Geodesic distances are calculated on the fly with the heat method [17] implemented in Geometry Central [68] and we use the Polyscope renderer [69] for visualization.
5 Experiments
We evaluate the performance of our surface-aware features in 3D correspondence matching, and in a downstream inference task of skinning weight regression.
5.1 Point-to-Point Correspondence Matching
@@
We follow the evaluation setup of Diff3F [24] and assess our surface-aware features in a correspondence matching task on human and animal shapes.
@@
We train a single autoencoder on a joint dataset consisting of 49 animal samples from the SMAL dataset [94] and 49 humans from the SURREAL dataset [35]. We choose 2 from each dataset for validation. For testing we use re-meshed versions of humans from SHREC’19 [53, 23] and animals from both SHREC’20 [25] and the animal-only subset of TOSCA [8].
@@
We compare our method against the unsupervised image-based Diff3F method [24], which also provides our base features, and against 3DCODED [35], DPC [44] and SE-OrNet [21], which have been trained on thousands of samples, while our method is trained on less than 100 samples.
@@
We report commonly used point correspondence metrics for $1024$ points per mesh [35, 44, 21, 24]111 We use the provided code and validate that we follow the same experimental procedure and metric definitions. . The correspondence error measures a distance between the computed correspondence point $\tau(\mathbf{p}_{n})$ (see Sec. 3) and the ground-truth correspondence point $\mathbf{t}_{n}^{gt}$ : $err=\frac{1}{n}\sum_{\mathbf{p}_{n}\in\mathcal{S}}\lVert\tau(\mathbf{p}_{n})-\mathbf{t}_{n}^{gt}\rVert_{2}^{2}$. The accuracy is the fraction of points with an error below a threshold $\epsilon\in[0,1]$: $acc(\epsilon)=\frac{1}{n}\sum_{\mathbf{p}_{n}\in\mathcal{S}}\mathbb{I}(\lVert\tau(\mathbf{p}_{n})-t_{n}^{gt}\rVert_{2}<\epsilon g)$, where $g$ is the maximal Euclidean distance in the target shape and $\mathbb{I}(.)$ is the indicator function.
@@ Results
We provide quantitative results in Tbl. 1 and qualitative comparisons in Fig. 5. We find that our method achieves the lowest error on SHREC’19 and TOSCA, despite being trained on fewer samples than the supervised baselines. Furthermore, we outperform the Diff3F base features on both SHREC datasets in terms of accuracy. While Diff3F achieves a higher accuracy at 1% threshold in TOSCA, Fig. 4 shows that the accuracy of our model is higher for thresholds above $\approx 2$%. This suggests that our method excels in the removal of outliers that can be caused by mismatched components.
Fig. 5 shows that Diff3F struggles to separate intraclass instances such as left and right legs. In contrast, the results confirm the effectiveness of our contrastive loss in mitigating this issue. We observe the same behavior for SHREC’20 in Fig. 3, which contains highly diverse animal shapes. Furthermore, our method generally produces visually smoother results (see Appendix 10.1 and the supplementary video).
@@
Table 1: Comparison of our 3D correspondence matching to prior works 3DC (3D-CODED [35]), DPC [44], SEN (SE-OrNet [21]), and Diff3F [24]. †) Numbers originate from [24], *) Experiments were replicated. Accuracy is for the commonly used $\epsilon=1$%. The per-column best results are bold and the second-to-best are underlined.
Figure 3: Qualitative comparison of correspondence matching on TOSCA (dense ground truth labels not available). Source shape (left) matched to target (right) using our and Diff3F features.
Figure 4: Correspondence accuracy ($\uparrow$ is better) at different error distance thresholds for our own and Diff3F features.
Figure 5: Qualitative Comparison on the SHREC’19 and TOSCA datasets with dense true correspondence labels. We show the source and target meshes with their ground truth correspondence labels (two left-most columns) in comparison to correspondences computed using our surface-aware features (forth column) and Diff3F base features (right-most column). We further highlight the correspondence error on the mesh surface (third and fifth column). The error colormap is normalized per sample by the maximal error over both methods to keep the error scale comparable across columns but not across rows. Our surface-aware features notably improve separation of the limb instances.
5.2 Exploration of Embedding Space
To illustrate the effect of our contrastive loss on feature separation and downstream applications, we compare the 2D projections of the DIFF3F base features with our surface-aware features, and we evaluate both as a base for the prediction of skinning weights.
Figure 6: Two shapes (left) and a PCA-based 2D projections of their aggregated Diff3F base features and our surface-aware features (right). Notice the separation of limbs in our result compared to Diff3F. Our features originate from the same encoder for both shapes.
Setup
We use our model, trained in Sec. 5.1, and create two dataset SMPLeval and SMALeval . The former consists of 50 randomly-sampled SMPL [8] shapes and poses from AMASS [52], while the latter consists of 50 randomly-sampled SMAL [94] shapes in canonical poses.
Embedding
We project Diff3F features aggregated from SMPLeval to two dimensions using principal component analysis (PCA). In Fig. 6, we visualize the projection for two selected shapes from the same dataset. We repeat this with our surface-aware features. To avoid bias, we derive the visualized colors from the true SMPL [48] skinning weights $\mathbf{w}_{n}\in\mathbb{R}^{B}$ for both methods, where $B$ is the skinning weight dimension. We repeat this process for SMALeval . In Fig. 6, we see that our method yields an interpretable embedding that separates the leg and hand instances for animals and humans despite not having access to extrinsic $(x,y,z)$ point positions. This remains a challenge for the Diff3F base features.
Downstream inference
Furthermore, we train a simple regressor to predict skinning weights of a kinematic model in a one-shot setup. The pointwise regressor $\mathcal{W}_{\mathbf{s}}(\mathbf{s}_{n})$ consists of a linear layer and a Softmax activation and regresses skinning weights $\bar{\mathbf{w}}_{n}$ from our surface-aware fetures or, in case of $\mathcal{W}_{\mathbf{f}}(\mathbf{f}_{n})$, from the base features. We train both models fivefold supervised with the Mean Squared Error (MSE) and true weights separately on the SMPLeval and SMALeval datasets, and we report test MSE for the remaining unused samples in the source datasets (see Fig. 7). Our features achieve lower errors and exhibit better robustness to instance ambiguities.
[Uncaptioned image] SMPLeval SMALeval Ours .0014±.0006 .0048±.0008 Diff3F .0058±.0024 .0086±.0028
Figure 7: Mean Squared Error of skinning weight regression ($\downarrow$ is better) and its distribution across SMPL mesh surface.
5.3 Ablations
We motivate our design choices by ablation of various parts of our method in Tbl. 2 following the setup of Sec. 5.1.

### v2

@@ 4.1. Setup
Unlike other approaches (Jiang et al., 2023; Deng et al., 2023), we do not require extrinsic canonical mesh alignment during training, because we rely only on intrinsic properties of individual meshes. Similar considerations were previously made for Functional maps (Ovsjanikov et al., 2012). Moreover, we inherit the rotation invariance of the base features demonstrated by Dutt et al. ([)Supplement “Robustness to Rotation”]dutt2024diffusion, although we observe a performance degradation if meshes are upside down and we avoid this in our inputs (see Appendix A.1).
Furthermore, our encoder is point-based and does not rely on shape completeness or a consistent topology. Both of these design choices favor generalization under transformations ranging from coordinate swap to shape reposing, and memory-unconstrained batch-based processing of even large shapes.
4.2. Separating Front Paw from Back Paw
Our embedding aims to separate multiple instances of the same class that are difficult to directly disambiguate in $\mathcal{F}_{\mathcal{M}}$. For example, consider the two surface points, $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, on the bear’s paws in Fig. 2. The prevalent semantic significance of the “paw” concept hinders the separability of the corresponding base features $\mathbf{f}_{1}$ and $\mathbf{f}_{2}$. Fig. 3 illustrates this for human arms and Diff3F features (Dutt et al., 2024). To solve this, we train a point-wise feature autoencoder, producing our surface-aware features $\mathcal{S}_{\mathcal{M}}\subset\mathbb{R}^{s}$ in its embedding space. We motivate the feasibility of separation by the prior observations that vision features additionally carry information about the global pose (Zhang et al., 2024a). We postulate that this enables our model to distinguish between part instances when guided by their intrinsic distance. We adopt the geodesic distance $d_{1,2}$ between $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ for this purpose. Following contrastive learning, we sample point pairs on a single shape to enforce $\phi(\mathbf{s}_{1},\mathbf{s}_{2})\approx d_{1,2}$ for $\mathbf{s}_{n}\in\mathcal{S}_{\mathcal{M}}$. We validate our choice of hyperspherical embedding space against Euclidean space in Sec. 5.5.
@@ 4.2. Separating Front Paw from Back Paw
We train a base feature encoder $\mathcal{E}(.)$, such that, following a normalization, $\mathbf{s}_{n}=\mathcal{E}(\mathbf{f}_{n})/\lVert\mathcal{E}(\mathbf{f}_{n})\lVert_{2}$ is a surface-aware feature $\mathbf{s}_{n}\in\mathbb{R}^{s}$ in a hypersphere embedding. During training, we randomly sample an unpaired training mesh $\mathcal{M}\in\mathbf{M}_{t}$ with base features $\mathcal{F}_{\mathcal{M}}$, which we encode pointwise to obtain $\mathcal{S}_{\mathcal{M}}$.
In each training iteration, we use furthest-point sampling to choose a random subset of $A$ anchor points $\mathbf{p}_{a}$ among the mesh vertices $\mathbf{p}_{i}\in\mathcal{M}{}$ and compute geodesic distances $d_{n,a}$ for each pair of a mesh and anchor point. We additionally rescale $d_{n,a}$ to a maximum of one, such that $d^{\prime}_{n,a}:=d_{n,a}/\max_{n,a}(d_{n,a})$, which removes the dependency on the scale of the mesh. We find this robust and leading to features later generalizing across morphologically equivalent shapes with different proportions (see an elephant vs. a giraffe in Fig. 8).
Subsequently, our contrastive loss preserves the rescaled geodesic distances in the embedding space:
(2) $\mathcal{L}_{c}=\frac{1}{NA}\sum_{n}^{N}\sum_{a}^{A}\left|d^{\prime}_{n,a}-\left(\frac{1-\phi(\mathbf{s}_{n},\mathbf{s}_{a})}{2}\right)\right|.$
This loss operates in a hyperspherical embedding space and utilizes cosine similarity mapped to the $[0,1]$ range. Hereby, $\mathcal{L}_{c}$ penalizes features close in the embedding space but distant on the shape surface and vice versa.
@@
(3) $\mathcal{L}_{r}=\frac{1}{N}\sum_{n}1-\phi(\mathbf{f}_{n},\mathbf{\bar{f}}_{n}).$
@@
Note that our training procedure, in contrast to related works (Lang et al., 2021; Deng et al., 2023; Chen et al., 2025), does not require target and source shape pairs.
4.3. Implementation
@@
During preprocessing, we rasterize our triangular meshes and precompute base features for all vertices following Sec. 3. We implement our autoencoder in PyTorch2 (Ansel et al., 2024) and use the Polyscope renderer (Sharp et al., 2019b) for visualizations. The encoder $\mathcal{E}$ is a Multilayer Perceptron (MLP) consisting of three blocks, where each block has two linear layers, SiLU activation (Elfwing et al., 2018), and layer normalization (Ba et al., 2016). The first layer in each block employs a skip connection (He et al., 2016), while the second reduces the dimensionality by a factor of two. With Diff3F (Dutt et al., 2024) as base features, $\mathcal{E}$ reduces feature dimensionality from $f=2048$ to $s=256$. The decoder $\mathcal{D}$ is a mirrored copy of the encoder. We train our model on NVIDIA RTX 3090 for 50k iterations with the AdamW optimizer (Loshchilov and Hutter, 2017) and a learning rate of $0.0001$ which takes $\approx 2$ hours.
We choose an exponential moving average (Polyak and Juditsky, 1992) of the model with the lowest validation loss, without the need for any correspondence labels. Geodesic distances for training are calculated on the fly with the heat method (Crane et al., 2017) implemented in Geometry Central (Sharp et al., 2019a). No geodesics are required during inference and the computational cost is determined by the Diff3F baseline with a only a negligible overhead from our shallow encoder $\mathcal{E}$. For a shape with 10k vertices, this is less than 5 milliseconds on top of $\approx 4$ minutes from Diff3F. Moreover, downstream tasks cost benefits from the smaller feature dimensionality. Functional maps are calculated with the base algorithm (Ovsjanikov et al., 2012), provided by the Diff3F implementation (Dutt et al., 2024).
5. Experiments
Here, we first motivate the benefits of our surface-aware feature embedding space by visualizing its distribution. Next, we evaluate their effectiveness in tasks with quantitative benchmarks including pose transfer, skinning weight regression and 3D correspondence matching. Finally, we analyze the impact of our design choices in an ablation experiment.
Training
We train a single autoencoder on a joint dataset consisting of 49 animal samples from the SMAL dataset (Zuffi et al., 2017) and 49 humans from the SURREAL dataset (Groueix et al., 2018). We choose 2 samples from each dataset for validation. We use this single shared model without any additional optimization for all experiments, unless stated otherwise.
5.1. Exploration of Embedding Space
To illustrate the effect of our contrastive loss on feature separation, we compare the 2D projections of the Diff3F (Dutt et al., 2024) base features with our surface-aware features.
Figure 3. Two shapes (left) and a PCA-based 2D projections of their aggregated Diff3F base features and our surface-aware features (right). Notice the separation of limbs in our result compared to Diff3F. Our features originate from the same encoder for both shapes. The animal legs appear merged along the sagittal plane due limitations of the PCA projection, but they remain disambiguated in our feature space as demonstrated in Fig. 9.
Setup
We create two dataset SMPLeval and SMALeval unseen during training. The former consists of 50 randomly-sampled SMPL (Loper et al., 2023) shapes and poses from AMASS (Mahmood et al., 2019), while the latter consists of 50 randomly-sampled SMAL (Zuffi et al., 2017; Biggs et al., 2018) shapes in canonical poses. For each sample, we obtain the base features and surface-aware features as described in Sec. 3 and Sec. 4
Embedding
We project Diff3F features aggregated from SMPLeval to two dimensions using principal component analysis (PCA). In Fig. 3, we visualize the projection for two selected shapes from the same dataset. We repeat this with our surface-aware features. To avoid bias, we derive the visualized colors from the true SMPL (Loper et al., 2023) skinning weights $\mathbf{w}_{n}\in\mathbb{R}^{B}$ for both methods, where $B$ is the skinning weight dimension. We repeat this process for SMALeval. In Fig. 3, our method yields an interpretable embedding that separates the leg and hand instances for animals and humans despite not having access to extrinsic $(x,y,z)$ point positions. This validates the suitability of our features for downstream tasks and highlights the limitations of the Diff3F base features.
5.2. One-shot Pose Transfer
Figure 4. One-shot pose transfer using our features, Diff3F features, or Geometric descriptors. MSE $\times 10^{-4}$ is reported for human shapes.
We evaluate the performance of the surface-aware features in a one-shot re-posing task for arbitrary 2-manifold meshes. We use our features to fit a Neural Jacobian Field (NJF) (Aigerman et al., 2022) between poses of two input shapes and then apply it to re-pose a new target shape as described in Appendix B.
We sample 5 input shape pairs from SMPLeval by choosing one challenging pose as a target and one random shape as an initial pose. The remaining shape samples serve as test inputs for pose transfer with known ground truth. We report MSE for 240 such test pairs (see upper row, Fig. 4). Next, we repeat the same procedure with the base features and with the Geometric descriptors (Geo) (Aigerman et al., 2022), consisting of the face centroid, face normals, and a Wave-Kernel Signature (Aubry et al., 2011). Finally, we provide additional qualitative results for transfer of animal poses from TOSCA to SMALeval in Fig. 4.
We observe that our surface-aware features outperform the base features both quantitatively and qualitatively by correctly distinguishing individual posed limbs. The Geometric descriptors also perform well for humans, but they struggle with larger input and output shape differences in animals. Our features perform well in both cases.
[Uncaptioned image] SMPLeval SMALeval Ours .0014±.0006 .0048±.0008 Diff3F .0058±.0024 .0086±.0028 Geo .0210±.0031 .0287±.0017
Figure 5. Mean Squared Error of skinning weight regression ($\downarrow$ is better) and its distribution across the SMPL mesh surface.
5.3. Skinning Weight Regression
We train a simple regressor to predict skinning weights of a kinematic model, based on a single training sample.
Our pointwise regressor $\mathcal{W}_{\mathbf{s}}(\mathbf{s}_{n})$ consists of a linear layer and a Softmax activation and regresses skinning weights $\bar{\mathbf{w}}_{n}$ from our surface-aware features or, in case of $\mathcal{W}_{\mathbf{f}}(\mathbf{f}_{n})$ and $\mathcal{W}_{\mathbf{g}}(\mathbf{g}_{n})$, from the base features or the Geometric descriptors, respectively. We train all models fivefold supervised with the Mean Squared Error (MSE) and true weights separately on the SMPLeval and dsmalours datasets, and we report test MSE for the remaining unseen samples in the source datasets (see Fig. 5). Our features achieve lower errors and exhibit better robustness to instance ambiguities than the two alternatives. It is worth noting that the relatively lower dimensionality of the Geometric descriptors and base features affects the number of regressor parameters, impact of which is not studied in this experiment. However, our conclusions also hold for a 2-layer MLP regressor with an equal hidden dimensionality of 106 which matches the dimensionality of the Geometric descriptors.
5.4. Point-to-Point Correspondence Matching
@@
Our features can be easily integrated into correspondence matching pipelines. Therefore, we replicate the evaluation setup of Diff3F (Dutt et al., 2024) and assess our surface-aware features in a correspondence matching task on human and animal shapes.
@@
For testing we use re-meshed versions of humans from SHREC’19 (Melzi et al., 2019; Donati et al., 2020) and animals from both SHREC’20 (Dyke et al., 2020) and the animal-only subset of TOSCA (Bronstein et al., 2008).
@@
We compare our method against the unsupervised image-based Diff3F method (Dutt et al., 2024), which also provides our base features, and against 3DCODED (Groueix et al., 2018), DPC (Lang et al., 2021) and SE-OrNet (Deng et al., 2023), which have been trained on thousands of samples, while our method is trained on less than 100 samples.
@@
We report commonly used point correspondence metrics for $1024$ points per mesh (Groueix et al., 2018; Lang et al., 2021; Deng et al., 2023; Dutt et al., 2024) [^1] We use the provided code and validate that we follow the same experimental procedure and metric definitions. . The correspondence error measures a distance between the computed correspondence point $\tau(\mathbf{p}_{n})$ (see Sec. 3) and the ground-truth correspondence point $\mathbf{t}_{n}^{gt}$ : $err=\frac{1}{n}\sum_{\mathbf{p}_{n}\in\mathcal{S}}\lVert\tau(\mathbf{p}_{n})-\mathbf{t}_{n}^{gt}\rVert_{2}^{2}$. The accuracy is the fraction of points with an error below a threshold $\epsilon\in[0,1]$: $acc(\epsilon)=\frac{1}{n}\sum_{\mathbf{p}_{n}\in\mathcal{S}}\mathbb{I}(\lVert\tau(\mathbf{p}_{n})-t_{n}^{gt}\rVert_{2}<\epsilon g)$, where $g$ is the maximal Euclidean distance in the target shape and $\mathbb{I}(.)$ is the indicator function.
@@ Results
We provide quantitative results in Tbl. 1 and qualitative comparisons in Fig. 8. We find that our method achieves the lowest error on SHREC’19 and TOSCA, despite being trained on fewer samples than the supervised baselines. Furthermore, we outperform the Diff3F base features on both SHREC datasets in terms of accuracy. While Diff3F achieves a higher accuracy at 1% threshold in TOSCA, Fig. 8 shows that the accuracy of our model is higher for thresholds above $\approx 2$%. This suggests that our method excels in the removal of outliers that can be caused by mismatched components. Additionally, adapting our features to Functional maps (Ovsjanikov et al., 2012) (+ FM) distributes the error towards a lower mean at the cost of accuracy, and maintains the beneficial comparison to Diff3F+FM.
Fig. 8 shows that Diff3F struggles to separate intraclass instances such as left and right legs. In contrast, the results confirm the effectiveness of our contrastive loss in mitigating this issue. We observe the same behavior for SHREC’20 in Fig. 8 (top), which contains highly diverse animal shapes. Furthermore, our method generally produces visually smoother results (see Appendix D.1 and the supplementary videos on the project website).
Table 1. Comparison of our 3D correspondence matching to prior works 3DC (3D-CODED (Groueix et al., 2018)), DPC (Lang et al., 2021), SEN (SE-OrNet (Deng et al., 2023)), and Diff3F (Dutt et al., 2024). †) Numbers originate from (Dutt et al., 2024), *) Experiments were replicated, x) Omitted due to non-manifold meshes, + FM) Semantic features combined with Functional Maps. Accuracy is for the commonly used $\epsilon=1$ %. The per-column best results are bold and the second-to-best results are underlined.
@@
Ours + FM err ↓ 0.24±0.64 x 3.54±7.59
acc ↑ 24.83±6.80 x 63.61±24.34
Figure 6. Qualitative comparison on the SHREC’19 and TOSCA datasets with dense true correspondence labels provided by their authors. We show the source and target meshes with their ground truth correspondence labels (the two left-most columns) in comparison to correspondences computed using our surface-aware features (the forth column) and Diff3F base features (the right-most column). We further highlight the correspondence error on the mesh surface (the third and the fifth column). The error colormap is normalized per sample by the maximal error over both methods to keep the error scale comparable across columns but not across rows. Our surface-aware features notably improve separation of the limb instances.
Figure 7. Correspondence accuracy ($\uparrow$ is better) at different error distance thresholds for our own and Diff3F features.
Figure 8. Qualitative comparison of correspondence matching on TOSCA and ShapeNet  (Chang et al., 2015) (dense ground truth labels not available). Source shape (left) matched to target (right) using our and Diff3F features.
Other Shapes
Our method is applicable beyond humanoid and animal shapes, which we show by training two additional encoders for a subset of 50 chairs and 50 airplanes from ShapeNet (Chang et al., 2015). Here, we uniformly resample the mesh vertices for a better surface coverage (Wang et al., 2022).
For unseen shapes in Fig. 8, our surface-aware features again better distinguish same-class instances such as chair legs and airplane wings, supporting a wider applicability of our methodology. More examples are shown in Appendix D.2 and the supplementary videos.
5.5. Ablations
We motivate our design choices by ablation on various parts of our method in Tbl. 2 following the setup of Sec. 5.4.